### v2

# Tree balance in phylogenetic models

Abstract
Tree shape statistics, particularly measures of tree (im)balance, play an important role in the analysis of the shape of phylogenetic trees. With applications ranging from testing evolutionary models to studying the impact of fertility inheritance and selection, or tumor development and language evolution, the assessment of tree balance is crucial. Currently, a multitude of at least 30 (im)balance indices can be found in the literature, alongside numerous other tree shape statistics.
This diversity prompts essential questions: How can we minimize the selection of indices to mitigate the challenges of multiple testing? Is there a preeminent balance index tailored to specific tasks? Previous studies comparing the statistical power of indices in detecting trees deviating from the Yule model have been limited in scope, utilizing only a subset of indices and alternative tree models.
This research expands upon the examination of index power, encompassing all established indices and a broader array of alternative models. Our investigation reveals distinct groups of balance indices better suited for different tree models, suggesting that decisions on balance index selection can be enhanced with prior knowledge. Furthermore, we present the R software package poweRbal which allows the inclusion of new indices and models, thus facilitating future research.
Keywords: tree balance, phylogenetic models, Yule model, macroevolution
1 Introduction
A key aspect in the study of evolution is to understand the forces that drive speciation and extinction processes and their effect on the evolution of species and taxa of higher level [16]. In the 1970s, the idea formed that such evolutionary processes can be modeled using stochastic processes (see, e.g., [39, 38, 45, 17, 37, 44]). Around the same time, researchers started assessing the shapes of (phylogenetic) trees via tree shape statistics (mostly tree balance indices, which measure the degree of imbalance/asymmetry in trees) as well as analyzing the probabilities of certain shape patterns under basic phylogenetic models [12, 18], and several biogeographic studies introduced the use of null models in macroevolution [41, 48, 47]. At the intersection of these fields, a whole research branch emerged, which is concerned with developing stochastic models that are good and useful explanations of (species) evolution based on the analysis of tree shapes (see, e.g., [43, 49, 29, 33, 34, 5, 7, 8, 52, 36, 41, 15, 1]). Intuitively, when the moments of a tree shape statistic under a given null model are known, or when the null distribution can be approximated through simulation, statistical methods to test a given tree against the null hypothesis (i.e., that the tree was obtained under the null model) can be devised. These results can then be used to decide whether a given null model is indeed a realistic model for evolution.
Over the last decades, a vast range of tree shape statistics on the one hand, and phylogenetic or macroevolutionary models on the other hand, have been developed. Due to rapid developments in both fields, studies combining both fields, i.e., investigating phylogenetic models through tree balance, have been limited in scope to a handful of tree balance indices and models (see, e.g., [20, 29, 21, 1, 7, 22, 25, 26]). As a result, the idea of using tree shape statistics in the study of phylogenetic models has not reached its full potential yet. Indeed, there has been no exhaustive analysis of the power of different tree shape statistics. In other words, there has been no exhaustive study of the question which tree shape statistics are best at detecting and testing specific evolutionary models.
The recent survey of tree balance indices by Fischer et al. [13] has provided the first pillar for this endeavor: a comprehensive list of all established tree balance indices and further tree shape statistics. These range from old and widely used indices such as the Sackin and Colless indices [42, 10], to fairly new approaches such as the rooted quartet index [11] or the symmetry nodes index [26]. As part of this manuscript, we provide the second pillar: an extensive overview and list of basic (phylogenetic) tree models and their parameters (see supplementary material [27]), which is more updated and comprehensive than earlier surveys (such as the excellent review by Mooers et al. [32]) on this topic.
The main purpose of our work, though, is to bridge these two pillars by providing the means to examine the power of all established tree shape statistics for a wide array of models. While we focus on their power to detect the Yule model against certain discrete-time models for the purpose of this paper, our general framework can easily be extended to other null and alternative models as well as new tree shape statistics. In particular, we introduce the R software package poweRbal, which can be used to perform analyses analogous to the ones presented here. This package currently includes all established tree balance indices as well as numerous macroevolutionary models, but is implemented in such a way that new tree shape statistics and new models can easily be incorporated. Detailed instructions on how to use the package are given in Section 5. In general, we suggest to use this software package as a precursor to subsequent analyses and research: The user specifies the research question (e.g., which null and alternative models to investigate), and inputs these choices into our software. Our software then provides the user with the most powerful tree shape statistics for this setting to be used in subsequent statistical analyses of the data. This minimizes problems related to multiple testing and ensures that no better (i.e., more powerful) tree shape statistic is overlooked.
Related work
As indicated above, our paper is not the first to analyze tree shape statistics in the context of evolutionary models, but previous studies have been limited in scope. Nonetheless, we briefly summarize these studies and highlight some alternative approaches that have been used in the literature. First, similar to our study, several authors have studied the power of tree shape statistics to detect deviations from a given phylogenetic null model (e.g., [20, 29, 21, 1, 7, 22, 25, 26]). Using the Yule model as the null model, these studies have employed a handful of tree shape statistics and a handful of alternative models in their investigation. While some tree shape statistics such as the Colless index [10] have been analyzed in several of these studies, others, such as the rooted quartet index [11], have not been included at all. However, we remark that very recently and independent of our study, Khurana et al. [28] used a large selection of the tree shape statistics categorized in [13] in a study aimed at discriminating between empirical trees and trees generated under a constant-rate birth-death model. Nevertheless, by including all established tree shape statistics as well as a large variety of alternative models in our analyses, we complement and expand the existing literature.
Next to analyzing the power of tree shape statistics in the context of macroevolutionary null models, tree shape statistics have also been used in testing for biogeographic null models (see, e.g., [22, p. 108]). Here, the aim is to decide whether a local subphylogeny is significantly more symmetric or asymmetric than expected if a random phylogeny of the same size is picked from the underlying larger tree.
Finally, there have been some approaches of evaluating tree shape statistics outside the realm of classic statistical hypotheses tests. For instance, Matsen [31] took a geometric approach to quantifying the power of tree shape statistics to differentiate between similar and different trees, where similarity of trees was measured using the so-called nearest neighbor interchange distance. Intuitively, the idea is that under good tree shape statistics, similar trees should receive similar values, whereas distant trees should receive different values. Hayati et al. [19] recently followed up on Matsen’s idea, albeit using a different resolution function (for a more detailed discussion of both approaches, see [19, pp. 5-10]).
2 Preliminaries
Before we can present our results, we introduce some general definitions and notation, where we mainly follow the terminology of [13]. Throughout this manuscript, $X$ denotes a non-empty finite set, which is often referred to as a taxon set or simply a species set. If not stated otherwise, we may assume $X=\{1,2,\ldots,n\}$.
Rooted binary phylogenetic trees and related concepts
[…]
Remark.
The leaves of a tree (sometimes also called “tips”) represent species or lineages (depending on the context), so we sometimes simplify explanations by saying, e.g., that “a leaf speciates” or “a leaf goes extinct”, which actually means that the species or lineage represented by that particular leaf undergoes said evolutionary event.
Tree balance indices
[…]
A very intuitive example of how to measure tree imbalance is given by the Colless index, which compares the number of leaves in the two maximal pending subtrees (which are formally defined in the supplementary material [27]) of each inner node and takes the sum over these absolute differences (see Figure 1 on the right).
In Tables 1 and 2 in the supplementary material [27] we list all tree (im)balance indices considered in this paper with a short description and their formal definition. We refer the reader to the respective references and [13] for further details.
(Phylogenetic) tree models
A probabilistic model of binary (phylogenetic) trees $P_{n}$ (short: (phylogenetic) tree model), with $n\geq 1$, is a family of probability mappings $P_{n}:\mathcal{BT}^{\ast}_{n}\rightarrow[0,1]$ or $P_{n}:\mathcal{BT}_{n}\rightarrow[0,1]$, respectively, associating a tree $T\in\mathcal{BT}^{\ast}_{n}$ or a phylogenetic tree $\mathcal{T}\in\mathcal{BT}_{n}$ with its probability under the model. For some models, explicit formulas for the probability of a particular (phylogenetic) tree are known, while for others this is not the case (yet). Many models implicitly describe the probability distribution by giving an algorithmic process that constructs a (phylogenetic) tree with $n$ leaves under the corresponding model.
Any phylogenetic tree model induces a tree model by setting the probability $P_{n}(T)$ of a rooted binary tree $T\in\mathcal{BT}^{\ast}_{n}$ as the sum of all probabilities $P_{n}(\mathcal{T})$ over all $\mathcal{T}\in\mathcal{BT}_{n}$ that have $T$ as their underlying topology. Since our aim is to evaluate tree balance indices that work on trees (and do not depend on leaf labels or edge lengths), we focus purely on the (induced) tree models.
A very foundational tree model is the Yule model [54], which is a pure birth model, where all species have the same constant rate of speciation at all times. To obtain a tree shape under the Yule model using a forward process, starting with a single leaf, a leaf is chosen uniformly at random and replaced by a cherry (see Figure 1 on the left). Finally, a phylogenetic tree under the Yule model is obtained by assigning leaf labels uniformly at random to a shape generated by this process. The Yule model plays a central role in phylogenetics and has been used for many decades. It is known under a multitude of names (Yule model, equal-rates-Markov model (ERM), random branching model, Markovian dichotomous branching model, Yule-Harding model, Yule-Harding-Kingman model (YHK), or simply Markovian model) of which we will only use Yule model in the remainder of this manuscript. While it can be described under various approaches, the probabilities to obtain a specific (phylogenetic) tree can be explicitly stated as follows (see, e.g., [51, Proposition 1] or [50, Proposition 3.2]):
$P_{Y,n}(\mathcal{T})=\frac{2^{n-1}}{n!}\cdot\prod\limits_{v\in\mathring{V}(T)}\frac{1}{n_{v}-1}\qquad\text{and}\qquad P_{Y,n}(T)=2^{n-1-s(T)}\cdot\prod\limits_{v\in\mathring{V}(T)}\frac{1}{n_{v}-1}.$
While the Yule model is one of the oldest and one of the most famous tree models, and it is often used as a null model in phylogenetic analyses, there are many more models available to describe different evolutionary settings. Section 3 in the supplementary material [27] provides an extensive collection of tree models discussed in the literature.
Figure 1: Left: Illustration of the Yule model using the forward process (“tree growing”). At each step from $n$ to $n+1$, a leaf is chosen uniformly at random and replaced by a cherry. Thus, under the Yule model, the caterpillar tree on four leaves, occurs with probability $P_{4}(T^{cat}_{4})=\frac{2}{3}$, whereas the fully balanced tree of height two occurs with probability $P_{4}(T^{fb}_{2})=\frac{1}{3}$.
Right: Calculation of the Colless index values, i.e., the sum of the absolute differences in the number of leaves in the two maximal pending subtrees of all inner nodes (marked in gray above the respective nodes), for $n=4$. As the Colless index is an imbalance index, more imbalanced trees are assigned higher values.
3 Methods: How to measure the power of balance indices
While there are various ways of comparing tree balance indices, in the present manuscript, we focus on comparing their power to distinguish alternative phylogenetic models from a null model $P^{0}_{n}$. As most previous studies (e.g., [20, 29, 21, 1, 7, 22, 25, 26]) we choose the Yule model as the null model. The Yule model also belongs to several parameterized tree model families, which allows us to observe the power of TSS as we gradually move away from the Yule model by changing the respective parameter(s). The null hypothesis is that the trees have been constructed under the null model.
We use two-tailed testing as a foundation for measuring the power of the tree balance indices. In particular, this implies that the alternative hypothesis is any deviation from the null model, i.e., that trees are more balanced or imbalanced than expected by chance. The level of confidence $\alpha$ is set to $5\%$, i.e., at most $5\%$ of trees constructed under the null model will be incorrectly assessed as non-$P^{0}_{n}$ (first type error) and the sample size $N_{d}$ is set to $10^{3}$. Then, given a set $TSS$ of tree shape statistics, a null model $P^{0}_{n}$, an alternative model $P^{a}_{n}$, and the number of leaves $n$, the testing procedure is as follows:
  1. (1)
    Approximate the null distributions of the TSS values: For each $t\in TSS$, sample $N_{d}$ many trees under the null model $P^{0}_{n}$ and compute their TSS values $v^{0}_{t}=(v^{0}_{(t,j)})_{j=1,...,N_{d}}$ (see Figure 2).
  2. (2)
    Compute the critical region: Using a level of significance of $\alpha=5\%$, compute the 0.025- as well as the 0.975-quantiles of the $v^{0}_{t}$ values for each $t$ yielding $A^{crit}_{t}=(-\infty,y_{0.025}^{*})\ \cup\ (y_{0.975}^{*},\infty)$ as the critical region (for trees with values in this region the null model is rejected) and $[y_{0.025}^{*},y_{0.975}^{*}]$ as the region of acceptance (the null model is maintained) for each $t$.
  3. (3)
    Test trees under alternative model: Construct $N_{d}$ many trees under $P^{a}_{n}$. For each tree $i$ compute the TSS values for each $t$ and set $x_{t,i}=1$ if the null model is rejected and $x_{t,i}=0$ otherwise.
  4. (4)
    Compute power: For each $t$ calculate the number $R_{t}$ of these trees for which the null model is rejected. Then, the power of $t$ is the proportion of rejected trees or the average over $x_{t}=(x_{t,i})_{i=1,...,N_{d}}$:
    $\text{power}_{(n,P^{0}_{n},P^{a}_{n})}(t)=\overline{x_{t}}=\frac{1}{N_{d}}\cdot\sum_{i=1}^{N_{d}}{x_{t,i}}=\frac{R_{t}}{N_{d}}.$
  5. (5)
    Compute confidence interval: Using the central limit theorem we know that for high $N_{d}$ we approximately have a normal distribution $\overline{x_{t}}\approx N\left(\mu,\frac{\sigma^{2}}{\sqrt{N_{d}}}\right)$ where $\mu$ is the overall expected value and $\sigma^{2}$ the (finite) variance of a random variable $x_{t,l}$. This allows us to estimate the radius of a $(1-\alpha)$%-confidence interval by approximating $\sigma$ with the sample standard deviation $sd_{x_{t}}$ (note that $z_{\alpha}\approx 1.96$ for $\alpha=5$%, i.e., a 95%-confidence interval):
    $\text{radius}_{(n,P^{0}_{n},P^{a}_{n})}(t)=z_{\alpha}\cdot\frac{\sigma}{\sqrt{N_{d}}}\approx 1.96\cdot\frac{\text{sd}_{x_{t}}}{\sqrt{N_{d}}}=1.96\cdot\frac{\sqrt{\frac{1}{N_{d}-1}\cdot\sum_{i=1}^{N_{d}}{(x_{t,i}-\overline{x_{t}})^{2}}}}{\sqrt{N_{d}}}=1.96\cdot\frac{\sqrt{\frac{R_{t}(N_{d}-R_{t})}{N_{d}(N_{d}-1)}}}{\sqrt{N_{d}}}.$
    By using $R_{t}(N_{d}-R_{t})\leq(N_{d}/2)^{2}$ we can calculate an upper limit for the radius, namely $1.96\cdot\sqrt{\frac{1}{4(N_{d}-1)}}$, which is approximately $0.031\approx 3\%$ for $N_{d}=10^{3}$.
Figure 2: Top: Histograms of the exact distributions of the Furnas rank and the Colless index under the Yule model, the PDA model, and the ETM for $n=16$. The Furnas rank is a balance index, i.e., trees with higher values can be considered as more balanced, and the Colless index is an imbalance index, i.e. trees with higher values can be considered as more imbalanced. The uniform distribution of the ETM on $\mathcal{BT}^{\ast}_{n}$ is directly visible as the Furnas rank is also a tree enumeration method, which assigns all trees in $\mathcal{BT}^{\ast}_{n}$ their unique number in $\{1,...,|\mathcal{BT}^{\ast}_{n}|\}$ (here $|\mathcal{BT}^{\ast}_{16}|=10,905$). The dotted lines indicate the 0.025- and the 0.975-quantile under the Yule model.
Bottom: Histograms of the distributions of the Colless index under the Aldous’ $\beta$ splitting model for several choices of $\beta$ (2,000 sampled trees for each $\beta$ in this case) for $n=30$.
Selection of TSS, null model, and alternative models
[…]
Second, we only consider suitable indices, which means that we had to exclude one index and limit the evaluation of another one to small $n$ for practical reasons: The Colijn-Plazotta ranking produces too high values and thus causes computational problems even for small $n$, like $n=14$, and therefore had to be omitted. Furthermore, due to its quadratic runtime, the Furnas rank has only been evaluated for $n=30$.
While the cherry index is not a tree balance index [13], it is included in our study as it has been used in similar contexts already [7, 31]. The complete list of TSS can be found in the legends of the graphics.
Furthermore, we use (representatives of) all discrete-time tree models $P_{n}$ listed in Section 3 in the supplementary material [27]. The only exceptions are the most simplistic models that only produce a single tree shape, e.g., the caterpillar or maximally balanced tree. The comparisons are done for $n=30$ and $100$ to cover some range of tree sizes.
The above-mentioned procedure is performed for all combinations of $n$, TSS, and alternative models.
In Section 5 we will illustrate how to do similar experiments using any set of tree shape statistics, any set of alternative models, and with regard to any null model with the help of our R software package poweRbal provided alongside this manuscript.
4 Results
This section contains the most interesting results, i.e., only a selection of alternative tree models, of the simulation studies described in Section 3. The complete collection of result figures can be found in Section 4 of the supplementary material [27].
In brief, all figures show the power (with 95%-confidence bands) of the various TSS to correctly identify trees generated under models distinct from the Yule model as not having been generated under the Yule model. The power is plotted either against the leaf number $n$ (Figure 3) or against the model’s parameters (Figures 49).
Comparing the different figures, several observations spring to mind. First, and maybe unsurprisingly, the curves look very different for different models. In some cases, there are few to no crossings between the power curves for different TSS (e.g., Figures 4 and 7), implying that the ranking of the TSS by their power does not change across the different settings tested, whereas for others, there are significant changes (e.g., Figures 3, 6, or 8). We also note that symmetric (parent species splits into two child species) versus asymmetric speciation (one child species splits off but parent species persists) significantly impact the results (see, e.g., Figure 8).
Second, there are some TSS that (almost) consistently outperform the others, such as the $\widehat{s}$-shape statistic, whereas others, such as the modified maximum difference in width, tend to perform rather poorly. In both cases, there are notable exceptions, though. In case of the $\widehat{s}$-shape statistic it is interesting to note that it is outperformed by other TSS (such as the $B_{2}$ index, the Colless index, the Sackin index, and the Mean $I^{\prime}$ index) for parameter-model combinations that are close to the Yule model (see Figures 6, 8). In case of the modified maximum difference in widths, exceptions to its overall relatively poor performance are for instance Aldous’ $\beta$ splitting model (Figure 4) and the biased speciation model (Figure 6). Moreover, while the power of TSS tends to increase with increasing $n$ (see Figure  3 and the differences between $n=30$ and $n=100$ in Figure 8), the power of the modified maximum difference in widths does not seem to follow this pattern.
Third, sometimes several TSS perform almost equally well or equally poorly (indicated by their curves being clumped together as it is for instance the case for the top three curves in Figure 7), whereas in other cases the curves are clearly separated (e.g., Figure 6). Interestingly, the total cophenetic index as well as the quadratic Colless index show extremely similar behavior (their lines are nearly always right on top of each other), and it would be interesting to investigate this further.
While there are many more observations to be made by carefully studying the different plots, these observations already show that there are significant performance differences between the various TSS, which need to be considered when TSS are used in the context of macroevolutionary models.
Figure 3: This figure displays the power of all TSS to correctly identify trees generated under the PDA model as not having been generated under the Yule model for several choices of $n$.
Figure 4: This figure displays the power of all TSS to correctly identify trees generated under the Aldous’ $\beta$ splitting model as not having been generated under the Yule model. Since for $\beta=0$, Aldous’ splitting model corresponds to the Yule model, all TSS rejected $\approx 0.05$ of the trees in that case (as specified with the level of significance).
Figure 5: This figure displays the power of all TSS to correctly identify trees generated under Ford’s $\alpha$ model as not having been generated under the Yule model. Since for $\alpha=0$, Ford’s $\alpha$ model corresponds to the Yule model, all TSS rejected $\approx 0.05$ of the trees in that case (as specified with the level of significance).
Figure 6: This figure displays the power of all TSS to correctly identify trees generated under the biased speciation model as not having been generated under the Yule model.
Figure 7: This figure displays the power of all TSS to correctly identify trees generated under the asymmetric simple Brownian model as not having been generated under the Yule model.
Figure 8: This figure displays the power of all TSS to correctly identify trees generated under the (a)symmetric punctuated(-intermittent) log-Brownian model as not having been generated under the Yule model. The starting trait value is $x_{0}=10$ with constant $\sigma_{x}=1$.
Figure 9: This figure displays the power of all TSS to correctly identify trees generated under the alternative birth-death model as not having been generated under the Yule model.
5 Software
We now present our software package poweRbal, implemented in the free programming language $\mathsf{R}$, which provides the readers with the means to do similar analyses as shown in the preceding section and adapt them to their needs. The package will shortly be published on CRAN. Using this package, the user can freely specify the set of tree shape statistics (all tree balance indices implemented in the package treebalance [13] are already included and can be used, but additional TSS can be added by the user as well), the null model, the set of alternative models (all tree models analyzed in this manuscript have already been included in the package and can be used, but additional models can be added by the user as well), the number of leaves $n$, and the method parameters like the sample size $N_{d}$ or the level of significance.
Below, we give a short introduction on how to use the package using illustrative examples. First, the package is loaded with library("poweRbal").
poweRbal provides the means to generate trees under various tree models. For example, if we want to generate a tree specifically under the PDA model, we can use genPDATree(n) (information on the function names of the individual tree models is given in the tables in Section 3 of the supplementary material [27]). Another way to access any tree model and produce any number of trees is to use the wrapper function genTrees(n, Ntrees, tm). With n we specify the number of leaves, with Ntrees the number of trees, and with tm the tree model either as a string/name if the tree model does not have or need a parameter or as a list of a string/name and parameter (see ?genTrees for more information on how to access the individual models). For example, we can use the following commands to generate 5 trees with 10 leaves each, once under the Yule model (default: rate 1), once under the PDA model, and once under Aldous’ $\beta$ splitting model with $\beta=-1$:
genTrees(n = 10, Ntrees = 5, tm = "yule")
genTrees(n = 10, Ntrees = 5, tm = "pda")
genTrees(n = 10, Ntrees = 5, tm = list("aldous", -1))
To compare the power of pre-implemented TSS under pre-implemented tree models, we can use the function powerCompare (see ?powerComp for detailed information). For this, we only have to specify the TSS, the alternative models, as well as the number of leaves. Optionally, the user can also specify the null model (by default the Yule model) and several other options that define the method (by default: sample size = 1,000, level of significance = 0.05, two-tailed test). For example, to use the Sackin and $B_{1}$ index (see tssInfo for information on how to access all available TSS) for the PDA and ETM models, we specify:
pc1 <- powerComp(tss = c("Sackin", "B1I"), alt_models = list("pda", "etm"), n = 10)
# Which yields the following matrix:
pda etm
Sackin 0.255 0.129
B1I 0.213 0.106
As another example, to use the Sackin and $B_{1}$ index for Aldous’ $\beta$ splitting model with $\beta=-1.5$, $\beta=-1$, and $\beta=-0.5$, we have:
pc2 <- powerComp(tss = c("Sackin", "B1I"), n = 10,
alt_models = list(list("aldous", -1.5), list("aldous", -1),
list("aldous", -0.5)))
# Which yields the following matrix:
Sackin 0.272 0.087 0.043
B1I 0.242 0.071 0.044
The first results can be plotted [^2] The figures in Section 4 have been created similarly, but with the help of the ggplot2 package for aesthetic reasons. with showPower(power_data = pc1) and the second with
showPower(power_data = pc2) or – better for representatives of parametrized tree model families as in this case – with showPower_param(power_data = pc2, model_param = c(-1.5, -1, -0.5), ylim = c(0,0.4)) (see the help files accessible with ?showPower and ?showPower_param for more information on customizing these plots).
To compare new TSS (the example below combines the Sackin index and the total $I^{\prime}$ index from [25, p.53] into a new TSS) with the existing ones, we can create a list object containing the function and a short name (and optionally more information as in tssInfo).
library("treebalance")
new_tss_func <- function(tree){
n <- length(tree\$tip.label)
return(sackinI(tree)+ 0.14*n*IbasedI(tree, method = "total",
correction = "prime", logs = F))
new_tss <- list(func = new_tss_func, short = "SackTotI")
Similarly, we can add new tree models by creating a list object containing the function which produces Ntrees-many trees with $n$ leaves. With that, it is also possible to use lists of trees, e.g., from a tree model that was implemented in a different programming language, as a basis for a null or alternative model. The ape package [35] provides means for importing common file formats like the Newick format. As an example for this, let treeList be a list of type multiphylo containing trees with the same number of leaves, say $n=30$. Then we could proceed as follows.
new_tm_func <- function(n , Ntrees = 1){
if (n == 30) {
selection <- sample.int(length(treeList), Ntrees)
return(treeList[selection])
} else {
stop("No such trees available.")
new_tm <- list(func = new_tm_func)
Then, we add the names of the list objects as follows. Here, the new tree model is used as an alternative model but it can be also used as the null model by setting null_model = "new_tm".
pc3 <- powerComp(tss = c("Sackin", "new_tss", "B1I"),
alt_models = list("pda", "new_tm"), n = 10)
showPower(power_data = pc3)
6 Conclusion and outlook
Tree shape statistics in general and tree balance indices in particular are powerful tools to find realistic models for the evolution of sets of species under investigation. However, not all these statistics are equally suitable for detecting every evolutionary model.
In most of our analyses, the $\widehat{s}$-shape statistic was the most powerful statistic discriminating alternative models from the Yule model. While this nicely complements a study by Blum at al. [8], who found that the $\widehat{s}$-shape statistic “warrants maximal power for rejecting the [Yule] model against the PDA [model]”, this is hardly surprising. It can easily be explained by the fact that the $\widehat{s}$-shape statistic was even introduced with that very purpose in mind. In fact, except for the normalizing constant of $P_{Y,n}$, the $\widehat{s}$-shape statistic corresponds to the logarithm of $P_{Y,n}$ (c.f. [46]), i.e., of the probability of generating a tree under the Yule model. However, while Blum et al. in their study explicitly did not claim that the $\widehat{s}$-shape statistic is “generally superior”, in our study we could see that it is indeed superior concerning many models, not only the PDA model. Some exceptions include the Brownian models, in which the $\widehat{s}$-shape statistic is outperformed by the $B_{2}$, Sackin and Colless indices, c.f. Figures 7 and 8. Khurana et al.  [28] also found that for their data, the $\widehat{s}$-shape statistic’s performance was mediocre, while the $B_{2}$ index, the variance of leaf depths, $I$-based indices, stairs2, and the normalized Colless index were most powerful in discriminating empirical trees from trees simulated under a constant-rate-birth-death model (which induces the same distribution on $\mathcal{BT}^{\ast}_{n}$ as the Yule model). Their findings concerning $B_{2}$ nicely complement a study by Bienvenu et al. [6], which used five different null models (one of which was the Yule model) and tested six different balance indices for their power (note that they did not include the $\widehat{s}$-shape statistic in their study, though). In this study, the authors found that the $B_{2}$-index is on average better than the other investigated indices. Note that in most of our studies, too, the $B_{2}$ index was amongst the most powerful indices.
As mentioned above, the choice of different null models can make a huge difference. However, even with the Yule model as a null model, there are huge performance differences between the different TSS. These differences might imply that there are different patterns of imbalance produced by the different models, some of which can only be recognized by a subset of tree shape statistics. Therefore, the choice of the most powerful TSS for each individual case is crucial.
However, while we have shown performance differences between various balance indices, most of them (interestingly, with the exception of maximum width over maximum depth) have in common that generally their power increases with increasing $n$, i.e., with an increasing number of species under investigation. This information might be helpful for biologists in order to plan their studies.
Due to limited space in this manuscript, we have only listed a few example comparisons here. However, a comprehensive list of a wider collection of TSS / model combinations induced by our manuscript can be found in the supplementary material [27].
We are confident that our manuscript will inspire future research. Concerning applications, our software package poweRbal is specifically designed in such a way that it also caters to researchers unfamiliar with the programming language R. Trees generated by other software tools can easily be loaded into our package (e.g., in the well-known Newick format) and used as a basis for an alternative model. Newly developed indices and models can easily be added to the package, so that the user can test several variations to see which one is most powerful in the given setting. poweRbal also enables researchers to perform a simulation study beforehand to decide on which TSS to use and to subsequently analyze their data with the respective optimal tools.
Concerning theoretical research, on the other hand, a goal for future developments could be to develop tools to guess the most likely model simply based on a given tree. First results in this regard were presented in [24], where the likelihood of pitchforks and cherries under Ford’s $\alpha$-model has been calculated, which enables users to find the most likely $\alpha$-value. However, as the possibilities of using pitchforks and cherries for estimating the best-fitting overall model are limited, more research in this regard is necessary.
Finally, our manuscript has focused on basic models, i.e., on models which contain clear instructions on how to generate (phylogenetic) trees. However, in the literature also more elaborate and complex models can be found (e.g., [32, pp. 9–13] and [30]). For instance, evolution can be modeled at the genome level, so that subsequently a phylogenetic tree can be reconstructed (e.g., [30]). For the latter, different reconstruction methods can be used, and it has been shown that these methods also have an impact on the resulting tree’s balance (see, e.g., [40, 23], [9, p. 1141])). Moreover, we have primarily focused on models from phylogenetics, but tree models also frequently occur in other areas such as population genetics, probability theory, or computer science. While some of the models used in these areas are closely related to the Yule model (see, e.g., [14] and Remark 3.1 in the supplement of the present manuscript), it would be an interesting direction of future research to include models such as Aldous’ continuum random tree model [2, 3, 4] in the analyses.
Acknowledgement.
MF and SK were supported by the project ArtIGROW, which is a part of the WIR!-Alliance “ArtIFARM – Artificial Intelligence in Farming”, and gratefully acknowledge the Bundesministerium für Bildung und Forschung (Federal Ministry of Education and Research, FKZ: 03WIR4805) for financial support. Moreover, we wish to thank Volkmar Liebscher for helpful advice on the development of the R-package and Annalena Franke for implementing first versions of several tree models.